# Travelling Waves in a Mathematical Model for Oncolytic Virotherapy

### v1

@@ 1 Introduction
In 1991, the first genetically engineered oncolytic virus (OV)—a thymidine kinase–deficient herpes simplex virus type 1 (HSV-1) mutant—was developed and used to treat malignant glioma in nude mice [viro_intro]. Oncolytic virotherapy (OVT) has since emerged as a promising cancer treatment strategy in which oncolytic viruses are introduced into a patient to selectively infect and eliminate tumor cells [LinDanni2023Ovbp].
Over the last two decades, mathematical models for oncolytic viruses were developed and analysed [WodarzKamorva009, KomorovaWodarz2010, Tian, eftimie, BhattDarshakKartikey2022Mtsd, storey, Jenner2019, MorselliDavid2023AaCM, arwa, alturqwi]. Besides clinically relevant conclusions on the administration of oncolytic virotherapy, it turned out that the basic oncolytic virus model has very interesting, and non-trivial, mathematical properties. The model is often formulated as a system of reaction-diffuion equations and it can lead to oscillations, spatial pattern formation, invasion fronts, spiral waves and spatio-temporal chaotic behavior [POOLADVAND2021108520, eftimie, Tejas2025]. The analysis of these patterns has mostly focussed on numerical simulations. In the present work, we rigorously establish the existence of travelling wave solutions for the basic oncolytic virus reaction–diffusion model, which describes the spatial interactions between uninfected cancer cells $C$, infected cancer cells $I$, and free virions $V$ on an unbounded spatial domain.
We notice that the three-component reaction-diffusion system is non-cooperative, hence the standard theory for monotone or cooperative systems (e.g., [FangJian2009MWfP, LiangXing2007Asos, MaShiwang2001TWfD, WeinbergerHansF.2002Aold, WuJianhong2001TWFo, WangHaiyan2011SSaT_epidemy_travekingwave]) cannot be applied. Instead, we construct appropriate upper and lower solutions and apply Schauder’s fixed point theorem to show existence of travelling waves with speeds in an interval $\bar{c},\infty)$.
@@ 1.1 Oncolytic Virotherapy Modeling
Early theoretical work on oncolytic virotherapy (OVT) was shaped by the models of Wodarz and Komarova [Woodarz2002, WodarzKamorva009, KomorovaWodarz2010], who established a general ODE framework capturing key features of tumor growth and viral spread. Their analyses revealed that OVT dynamics fall into fast- and slow-spread regimes, and that treatment outcomes depend sensitively on assumptions about infection kinetics. These studies provided a foundation for later spatial and PDE-based models of tumor–virus interactions. Building on these ideas, Tian [Tian] introduced a three-variable ordinary differential system tracking healthy cancer cells, infected cells, and free virus, showing how viral burst size can determine therapeutic success. More detailed spatial models soon followed. Pooladvand et al.[POOLADVAND2021108520] developed a three-dimensional, radially symmetric PDE model for adenovirus therapy in solid tumors, allowing susceptible cells, infected cells, and free virus to diffuse. Their analysis showed that increasing viral infectivity alone does not guarantee tumor elimination—consistent with experimental observations that virotherapy rarely succeeds as a standalone treatment. Their work helped motivate further spatial modelling, including the diffusion-based system of [arwa], which examined temporal and spatial infection patterns in the absence of immune effects. Additional computational studies have explored why OVT can fail. Bhatt et al. [BhattDarshakKartikey2022Mtsd] combined immersed-boundary and Voronoi-based simulations to identify key limiting mechanisms: rapid loss of infected cells (leading to premature viral clearance), emergence of virus-resistant tumor cells, and insufficient viral spread rates. Morselli et al. [MorselliDavid2023AaCM] used a stochastic agent-based model to examine how spatial constraints and microenvironmental factors shape viral propagation, comparing random versus pressure-driven viral movement and reproducing infection patterns seen in earlier biological and theoretical studies. Although these agent-based and high-dimensional spatial models provide rich biological insight, their computational cost often limits extensive analytical investigation. For this reason, reaction–diffusion models remain valuable for uncovering the spatial structure of viral spread and identifying conditions favoring tumor control or eradication. Baadbulla et al. [arwa] extend the framework of Pooladvand et al. [POOLADVAND2021108520]. Pooladvands original model—formulated in 3-D radial symmetry—displayed transitions to coexistence states and oscillatory dynamics through a Hopf bifurcation, matching oscillations previously reported in experimental systems. Baabdulla et al [arwa] removed the radial symmetry and analyzed the full 2-D and 3-D diffusion problem. They computed invasion speeds numerically, and compared spatial infection profiles. They found complex spatiotemporal behaviours that are often seen in excitable media [Kuramoto, Sandstede]. Their analysis highlighted how spatial heterogeneity shapes viral spread and helps clarify the circumstances under which OVT can suppress—or nearly eradicate—tumor mass.
@@
In the absence of oncolytic viruses (OVs), tumors typically develop an immunosuppressive microenvironment that suppresses the host immune response. The introduction of OVs, however, induces pro-inflammatory signalling within the tumor and can substantially alter tumor–immune dynamics. Motivated by the work of Al-Tuwairqi et al. [alturqwi], a reaction–diffusion model was proposed in [NegarThomas1] to capture OVT–immune system interactions. This model extends that of Baabdulla et al. [arwa] by incorporating a fourth variable $Y$ representing the immune response. Their analysis demonstrated the existence of stacked travelling waves and, through the computation of associated wave speeds, indicated that OVT as a monotherapy is unlikely to achieve complete tumor eradication. This was further explored in [NegarTHomas2], where strategies to enhance the efficacy of OVT were investigated, with particular emphasis on the importance of combination treatments involving immunotherapies. These studies highlight that explicitly modelling the immune response is essential for understanding the full therapeutic potential of OVT.
@@
Eftimie et al. [eftimie] developed an ODE model incorporating both lymphoid and peripheral immune compartments to describe the interactions among uninfected and infected tumor cells, memory and effector immune cells, and two viral types (Ad and VSV). Their model successfully reproduced experimental tumor growth and immune-response patterns, identifying conditions under which sustained tumor elimination is possible. Importantly, they showed that viral persistence alone cannot clear the tumor without a coordinated anti-tumor immune response.
@@
Storey et al. [storey] further distinguished the roles of innate and adaptive immunity in OVT, noting that the innate response may act too rapidly and clear the virus before sufficient tumor infection occurs. To overcome this limitation, combination therapies pairing OVT with PD-1/PD-L1 checkpoint inhibition have been explored. Such approaches mitigate T-cell exhaustion, enhance immune-mediated tumor clearance, and reduce the viral infectivity threshold required for therapeutic success. Their findings also underscore the importance of treatment scheduling: administering a second viral dose too early can redirect the immune response toward viral, rather than tumor, elimination, thereby reducing overall therapeutic efficacy.
@@ 1.2 Travelling Waves
Modeling with partial differential equations (PDEs)—and reaction–diffusion equations in particular—has had a profound influence on the mathematical study of biological populations, with significant applications in cell biology [SherrattJA1992Ewht_murray_cell_traveling, LagergrenJohnH.2020Bnng_Cell_travelingwave], ecology [SkellamJ.G.1991Rdit_ecology_travelingwave, LevinSimonA.2003TEAE_ecology_travelingwave], and epidemiology [ZhangTianran2014Eotw_wang_epidomy_travelingwave, WangHaiyan2011SSaT_epidemy_travekingwave]. Many works (e.g., [murray, Britton, volpretVolpretVolpret, smoller_alma991043704819909116]) have investigated travelling waves and spatial dynamics for $n$-dimensional $(n\geq 1)$ reaction–diffusion systems of the form
@@
for $1\leq i\leq n$. Here $u_{i}(x,t)$ denotes the density of species of type $i=1,\dots,n$, $D_{i}$ denotes their diffusion coefficient, and $f_{i}(u_{1},\dots,u_{n})$ describes their birth, death, and other interactions. Such models give rise to travelling–wave solutions that represent invasion fronts. The most prominent model arises for one species $n=1$ and is known as the Fisher–KPP model [murray, Britton]. It is a classical and widely studied prototype for biological invasions and describes the spatiotemporal dynamics of a population whose individuals migrate via linear diffusion and proliferate according to logistic-type growth.
In cancer biology, reaction–diffusion frameworks have been especially useful in describing tumor invasion, particularly when the spread of malignant cells is coupled to degradation of surrounding tissue [El-HachemMaud2021Twao_travelingwave].
The system (1) is called cooperative if each $f_{i}(u)$ is non-decreasing in all components of $u$, with the possible exception of the $i$th one. Cooperative systems have been extensively studied, and a general travelling–wave theory is available in this setting [FangJian2009MWfP, WeinbergerHansF.2002Aold, LiangXing2007Asos].
For non-cooperative systems—characterized by kinetic Jacobians that include negative off-diagonal entries—this general theory does not apply. The model studied in this work (system (3)) is non-cooperative and the above mentioned theory is not available. Instead we follow the framework of [ThiessenRyan2026Twst] to construct upper and lower solutions that connect correctly to the left boundary and remain uniformly bounded throughout the domain. This framework then enables the application of Schauder’s fixed-point theorem, yielding the existence of travelling–wave solution for speeds in an interval $\bar{c},\infty)$.
@@ 2 The mathematical model
In this section, we will study the A. Baabdulla [arwa] reaction diffusion model on a rectangular or smooth domain $\Omega\subset\mathbb{R}^{n}$.
The proposed model from [arwa] is given by
@@
$\hat{C}=\frac{C}{L},\quad\hat{I}=\frac{I}{L},\quad\hat{V}=\frac{\beta V}{r},\quad\hat{t}=\tau t,\quad\hat{x}=x\sqrt{\frac{\tau}{D_{V}}}\ \text{, and}\quad\hat{D}=\frac{D}{D_{V}}.$
@@ 2.1 Analysis of the model
The analysis of the equilibrium points $E_{0}$, $E_{1}$, and $E_{2}$ was carried out in detail in [arwa]. We show a bifurcation diagram in Figure 1 with $\theta$ as bifurcation parameter. For low values of $\theta$, the cancer only state $E_{1}$ is stable, and by increasing $\theta$, it undergoes a transcritical bifurcation at $\theta_{t}=\gamma a$, and the coexistence state, $E_{2}$, becomes biologically relevant and stable. Upon further increase of $\theta$, we obtain a Hopf bifurcation at $\theta_{H}$ with periodic orbits arising for $\theta>\theta_{H}$. For our further analysis, we assume the coexistence equilibrium exists, hence we assume throughout:
@@ 3.1 The travelling wave problem
The system (3.1) is a second-order inhomogeneous ODE and we find an alternative formulation using the method of variation of parameters. This transformation has shown to be useful for travelling wave analysis in the work of Fang and Zhao [FangJian2009MWfP]. We define $U\coloneqq(B,I,V)$, matrix $\mathcal{D}$, and the function $f$ as
@@ 3.2 Dispersion relation for the invasion wave speed
The eigenvalue $\lambda_{1}(\rho)$ in (3.2) has the associated eigenvector $\nu_{1}=(1,0,0)$, which corresponds to a direction purely in the cancer component: it involves neither infected cells nor virus. In fact, $\lambda_{1}$ is determined entirely by the BB–equation and represents the spatial decay rate of the stable manifold of the classical Fisher–KPP equation at the equilibrium $C=1$ (equivalently, $B=0$). It is well known that this mode cannot generate a bounded travelling wave solution, and we therefore exclude this case from further consideration (see, for example, [Britton, murray]).
@@ 3.3 Conditions for the construction of upper and lower solutions
In Figure 2, we present four different cases illustrating the relative alignment of the critical points $\rho_{m}$, $\rho_{*}$, $\tilde{\rho}$, and $\check{\rho}$. Figures 2(a) and 2(c) display the curves $S(\rho)$, $H_{1}(\rho)$, $H_{2}(\rho)$, $H_{3}(\rho)$, and $H_{4}(\rho)$ in the parameter regime $a>1$. In Figure 2(a), since $a>1$, the curve $H_{1}(\rho)$ lies above $H_{2}(\rho)$. However, there is no intersection between $H_{1}(\rho)$ and $S(\rho)$ for $\rho<\rho_{m}$. Similarly, $H_{3}(\rho)$ and $S(\rho)$ do not intersect for $\rho<\rho_{m}$. The curve $H_{4}(\rho)$ intersects $S(\rho)$ for $\rho>\rho_{m}$, and the point $(\rho_{m},c_{m})$ lies within the shaded region. Consequently, in this case we have $\bar{\rho}=\rho_{m}$. The value of $\gamma$ used in Figures 2(a) and 2(c) is taken from [arwa] and equals $\gamma=40/3$. By contrast, Friedman et al. [FriedmanAvner2006GVEo] reported a viral clearance rate of $\gamma=1.25$. This smaller value of $\gamma$ yields a less restrictive shaded region, as illustrated in Figures 2(b) and 2(d). Figure 2(b) depicts a possible, though biologically unrealistic, parameter regime in which the curve $H_{1}(\rho)$ intersects $S(\rho)$, as does $H_{3}(\rho)$. In this situation, the smallest of these intersection points lying within the shaded region determines $\bar{\rho}$.
@@ 4.1 Existence of travelling wave solution for $c\geq\bar{c}$
Thus, by Theorem 3.4.2 in [Hillen_functional], $T$ is also a compact operator. ∎
Combining the results of the previous lemmas, we apply Schauder’s Fixed Point Theorem [Hillen_functional] to establish the existence of a travelling wave solution for our system of equations (3.1).
@@
Step 2:For $c=\bar{c}$, we employ the limiting argument of Fang and Zhao [FangJian2009MWfP] (Theorem 3.1). Choose a sequence $\{c_{n}\}_{n\geq 1}\subset(\bar{c},\infty)$ such that

### v2

@@ 1 Introduction
In 1991, the first genetically engineered oncolytic virus (OV)—a thymidine kinase–deficient herpes simplex virus type 1 (HSV-1) mutant—was developed and used to treat malignant glioma in nude mice [19]. Oncolytic virotherapy (OVT) has since emerged as a promising cancer treatment strategy in which oncolytic viruses are introduced into a patient to selectively infect and eliminate tumor cells [17].
Over the last two decades, mathematical models for oncolytic viruses were developed and analysed [35, 12, 31, 6, 4, 29, 11, 22, 2, 1]. Besides clinically relevant conclusions on the administration of oncolytic virotherapy, it turned out that the basic oncolytic virus model has very interesting, and non-trivial, mathematical properties. The model is often formulated as a system of reaction-diffuion equations and it can lead to oscillations, spatial pattern formation, invasion fronts, spiral waves and spatio-temporal chaotic behavior [24, 6, 3]. The analysis of these patterns has mostly focussed on numerical simulations. In the present work, we rigorously establish the existence of travelling wave solutions for the basic oncolytic virus reaction–diffusion model, which describes the spatial interactions between uninfected cancer cells $C$, infected cancer cells $I$, and free virions $V$ on an unbounded spatial domain.
We notice that the three-component reaction-diffusion system is non-cooperative, hence the standard theory for monotone or cooperative systems (e.g., [8, 16, 18, 34, 37, 33]) cannot be applied. Instead, we construct appropriate upper and lower solutions and apply Schauder’s fixed point theorem to show existence of travelling waves with speeds in an interval $\bar{c},\infty)$.
@@ 1.1 Oncolytic Virotherapy Modeling
Early theoretical work on oncolytic virotherapy (OVT) was shaped by the models of Wodarz and Komarova [36, 35, 12], who established a general ODE framework capturing key features of tumor growth and viral spread. Their analyses revealed that OVT dynamics fall into fast- and slow-spread regimes, and that treatment outcomes depend sensitively on assumptions about infection kinetics. These studies provided a foundation for later spatial and PDE-based models of tumor–virus interactions. Building on these ideas, Tian [31] introduced a three-variable ordinary differential system tracking healthy cancer cells, infected cells, and free virus, showing how viral burst size can determine therapeutic success. More detailed spatial models soon followed. Pooladvand et al.[24] developed a three-dimensional, radially symmetric PDE model for adenovirus therapy in solid tumors, allowing susceptible cells, infected cells, and free virus to diffuse. Their analysis showed that increasing viral infectivity alone does not guarantee tumor elimination—consistent with experimental observations that virotherapy rarely succeeds as a standalone treatment. Their work helped motivate further spatial modelling, including the diffusion-based system of [2], which examined temporal and spatial infection patterns in the absence of immune effects. Additional computational studies have explored why OVT can fail. Bhatt et al. [4] combined immersed-boundary and Voronoi-based simulations to identify key limiting mechanisms: rapid loss of infected cells (leading to premature viral clearance), emergence of virus-resistant tumor cells, and insufficient viral spread rates. Morselli et al. [22] used a stochastic agent-based model to examine how spatial constraints and microenvironmental factors shape viral propagation, comparing random versus pressure-driven viral movement and reproducing infection patterns seen in earlier biological and theoretical studies. Although these agent-based and high-dimensional spatial models provide rich biological insight, their computational cost often limits extensive analytical investigation. For this reason, reaction–diffusion models remain valuable for uncovering the spatial structure of viral spread and identifying conditions favoring tumor control or eradication. Baadbulla et al. [2] extend the framework of Pooladvand et al. [24]. Pooladvands original model—formulated in 3-D radial symmetry—displayed transitions to coexistence states and oscillatory dynamics through a Hopf bifurcation, matching oscillations previously reported in experimental systems. Baabdulla et al [2] removed the radial symmetry and analyzed the full 2-D and 3-D diffusion problem. They computed invasion speeds numerically, and compared spatial infection profiles. They found complex spatiotemporal behaviours that are often seen in excitable media [13, 25]. Their analysis highlighted how spatial heterogeneity shapes viral spread and helps clarify the circumstances under which OVT can suppress—or nearly eradicate—tumor mass.
@@
In the absence of oncolytic viruses (OVs), tumors typically develop an immunosuppressive microenvironment that suppresses the host immune response. The introduction of OVs, however, induces pro-inflammatory signalling within the tumor and can substantially alter tumor–immune dynamics. Motivated by the work of Al-Tuwairqi et al. [1], a reaction–diffusion model was proposed in [21] to capture OVT–immune system interactions. This model extends that of Baabdulla et al. [2] by incorporating a fourth variable $Y$ representing the immune response. Their analysis demonstrated the existence of stacked travelling waves and, through the computation of associated wave speeds, indicated that OVT as a monotherapy is unlikely to achieve complete tumor eradication. This was further explored in [20], where strategies to enhance the efficacy of OVT were investigated, with particular emphasis on the importance of combination treatments involving immunotherapies. These studies highlight that explicitly modelling the immune response is essential for understanding the full therapeutic potential of OVT.
@@
Eftimie et al. [6] developed an ODE model incorporating both lymphoid and peripheral immune compartments to describe the interactions among uninfected and infected tumor cells, memory and effector immune cells, and two viral types (Ad and VSV). Their model successfully reproduced experimental tumor growth and immune-response patterns, identifying conditions under which sustained tumor elimination is possible. Importantly, they showed that viral persistence alone cannot clear the tumor without a coordinated anti-tumor immune response.
@@
Storey et al. [29] further distinguished the roles of innate and adaptive immunity in OVT, noting that the innate response may act too rapidly and clear the virus before sufficient tumor infection occurs. To overcome this limitation, combination therapies pairing OVT with PD-1/PD-L1 checkpoint inhibition have been explored. Such approaches mitigate T-cell exhaustion, enhance immune-mediated tumor clearance, and reduce the viral infectivity threshold required for therapeutic success. Their findings also underscore the importance of treatment scheduling: administering a second viral dose too early can redirect the immune response toward viral, rather than tumor, elimination, thereby reducing overall therapeutic efficacy.
@@ 1.2 Travelling Waves
Modeling with partial differential equations (PDEs)—and reaction–diffusion equations in particular—has had a profound influence on the mathematical study of biological populations, with significant applications in cell biology [26, 14], ecology [27, 15], and epidemiology [38, 33]. Many works (e.g., [23, 5, 32, 28]) have investigated travelling waves and spatial dynamics for $n$-dimensional $(n\geq 1)$ reaction–diffusion systems of the form
@@
for $1\leq i\leq n$. Here $u_{i}(x,t)$ denotes the density of species of type $i=1,\dots,n$, $D_{i}$ denotes their diffusion coefficient, and $f_{i}(u_{1},\dots,u_{n})$ describes their birth, death, and other interactions. Such models give rise to travelling–wave solutions that represent invasion fronts. The most prominent model arises for one species $n=1$ and is known as the Fisher–KPP model [23, 5]. It is a classical and widely studied prototype for biological invasions and describes the spatiotemporal dynamics of a population whose individuals migrate via linear diffusion and proliferate according to logistic-type growth.
In cancer biology, reaction–diffusion frameworks have been especially useful in describing tumor invasion, particularly when the spread of malignant cells is coupled to degradation of surrounding tissue [7].
The system (1) is called cooperative if each $f_{i}(u)$ is non-decreasing in all components of $u$, with the possible exception of the $i$th one. Cooperative systems have been extensively studied, and a general travelling–wave theory is available in this setting [8, 34, 16].
For non-cooperative systems—characterized by kinetic Jacobians that include negative off-diagonal entries—this general theory does not apply. The model studied in this work (system (3)) is non-cooperative and the above mentioned theory is not available. Instead we follow the framework of [30] to construct upper and lower solutions that connect correctly to the left boundary and remain uniformly bounded throughout the domain. This framework then enables the application of Schauder’s fixed-point theorem, yielding the existence of travelling–wave solution for speeds in an interval $\bar{c},\infty)$.
@@ 2 The mathematical model
In this section, we will study the A. Baabdulla [2] reaction diffusion model on a rectangular or smooth domain $\Omega\subset\mathbb{R}^{n}$.
The proposed model from [2] is given by
@@
$\hat{C}=\frac{C}{L},\quad\hat{I}=\frac{I}{L},\quad\hat{V}=\frac{\beta V}{r},\quad\hat{t}=rt,\quad\hat{x}=x\sqrt{\frac{r}{D_{V}}}\ \text{, and}\quad\hat{D}=\frac{D}{D_{V}}.$
@@ 2.1 Analysis of the model
The analysis of the equilibrium points $E_{0}$, $E_{1}$, and $E_{2}$ was carried out in detail in [2]. We show a bifurcation diagram in Figure 1 with $\theta$ as bifurcation parameter. For low values of $\theta$, the cancer only state $E_{1}$ is stable, and by increasing $\theta$, it undergoes a transcritical bifurcation at $\theta_{t}=\gamma a$, and the coexistence state, $E_{2}$, becomes biologically relevant and stable. Upon further increase of $\theta$, we obtain a Hopf bifurcation at $\theta_{H}$ with periodic orbits arising for $\theta>\theta_{H}$. For our further analysis, we assume the coexistence equilibrium exists, hence we assume throughout:
@@ 3.1 The travelling wave problem
The system (3.1) is a second-order inhomogeneous ODE and we find an alternative formulation using the method of variation of parameters. This transformation has shown to be useful for travelling wave analysis in the work of Fang and Zhao [8]. We define $U\coloneqq(B,I,V)$, matrix $\mathcal{D}$, and the function $f$ as
@@ 3.2 Dispersion relation for the invasion wave speed
The eigenvalue $\lambda_{1}(\rho)$ in (3.2) has the associated eigenvector $\nu_{1}=(1,0,0)$, which corresponds to a direction purely in the cancer component: it involves neither infected cells nor virus. In fact, $\lambda_{1}$ is determined entirely by the BB–equation and represents the spatial decay rate of the stable manifold of the classical Fisher–KPP equation at the equilibrium $C=1$ (equivalently, $B=0$). It is well known that this mode cannot generate a bounded travelling wave solution, and we therefore exclude this case from further consideration (see, for example, [5, 23]).
@@ 3.3 Conditions for the construction of upper and lower solutions
In Figure 2, we present four different cases illustrating the relative alignment of the critical points $\rho_{m}$, $\rho_{*}$, $\tilde{\rho}$, and $\check{\rho}$. Figures 2(a) and 2(c) display the curves $S(\rho)$, $H_{1}(\rho)$, $H_{2}(\rho)$, $H_{3}(\rho)$, and $H_{4}(\rho)$ in the parameter regime $a>1$. In Figure 2(a), since $a>1$, the curve $H_{1}(\rho)$ lies above $H_{2}(\rho)$. However, there is no intersection between $H_{1}(\rho)$ and $S(\rho)$ for $\rho<\rho_{m}$. Similarly, $H_{3}(\rho)$ and $S(\rho)$ do not intersect for $\rho<\rho_{m}$. The curve $H_{4}(\rho)$ intersects $S(\rho)$ for $\rho>\rho_{m}$, and the point $(\rho_{m},c_{m})$ lies within the shaded region. Consequently, in this case we have $\bar{\rho}=\rho_{m}$. The value of $\gamma$ used in Figures 2(a) and 2(c) is taken from [2] and equals $\gamma=40/3$. By contrast, Friedman et al. [9] reported a viral clearance rate of $\gamma=1.25$. This smaller value of $\gamma$ yields a less restrictive shaded region, as illustrated in Figures 2(b) and 2(d). Figure 2(b) depicts a possible, though biologically unrealistic, parameter regime in which the curve $H_{1}(\rho)$ intersects $S(\rho)$, as does $H_{3}(\rho)$. In this situation, the smallest of these intersection points lying within the shaded region determines $\bar{\rho}$.
@@ 4.1 Existence of travelling wave solution for $c\geq\bar{c}$
Thus, by Theorem 3.4.2 in [10], $T$ is also a compact operator. ∎
Combining the results of the previous lemmas, we apply Schauder’s Fixed Point Theorem [10] to establish the existence of a travelling wave solution for our system of equations (3.1).
@@
Step 2:For $c=\bar{c}$, we employ the limiting argument of Fang and Zhao [8] (Theorem 3.1). Choose a sequence $\{c_{n}\}_{n\geq 1}\subset(\bar{c},\infty)$ such that